# Continual Imitation Learning for Prosthetic Limbs

### v1

@@ Abstract
Lower limb amputations and neuromuscular impairments severely restrict mobility, necessitating advancements beyond conventional prosthetics. Motorized bionic limbs offer promise, but their utility depends on mimicking the evolving synergy of human movement in various settings. In this context, we present a novel model for bionic prostheses’ application that leverages camera-based motion capture and wearable sensor data, to learn the synergistic coupling of the lower limbs during human locomotion, empowering it to infer the kinematic behavior of a missing lower limb across varied tasks, such as climbing inclines and stairs. We propose a model that can multitask, adapt continually, anticipate movements, and refine. The core of our method lies in an approach which we call “multitask prospective rehearsal” that anticipates and synthesizes future movements based on the previous prediction and employs a corrective mechanism for subsequent predictions. We design an evolving architecture that merges lightweight, task-specific modules on a shared backbone, ensuring both specificity and scalability. We empirically validate our model against various baselines using real-world human gait datasets, including experiments with transtibial amputees, which encompass a broad spectrum of locomotion tasks. The results show that our approach consistently outperforms baseline models, particularly under scenarios affected by distributional shifts, adversarial perturbations, and noise.
I Introduction
@@
Lower limb amputation and neuromuscular disorders detrimentally affect natural locomotion, compromising individuals’ quality of life [1]. Those afflicted often depend on assistive technologies like prosthetics or orthoses to regain daily mobility [1]. However, conventional passive prosthetics often fall short in replicating natural gait during diverse activities, from basic transitions like standing from a seated position to more dynamic actions like running or navigating slopes [1]. The advent of powered prosthetics, equipped with integrated motors, promises a more naturalistic gait. Nonetheless, this advancement requires a model capable of effectively approximating the complexities of human gait synergy to accurately estimate the necessary motor commands. While finite-state machines are adequate for rudimentary scenarios [2, 3], their construction becomes infeasible for finer gait phase resolution or multiple locomotion tasks [4]. A promising alternative involves leveraging models trained on able-bodied human demonstrations to intuitively infer amputee limb motion. This methodology not only replicates gait’s inherent fluidity but also facilitates seamless transitions across varying gaits, eliminating the need for rigid rules or heuristics. The learned gait models could be personalized by training on a wide variety of human demonstrations, such that references with similar anthropomorphic features as that of the physically impaired subject are included.
@@
Recent studies attest to the efficacy of learning-based models in gait and prosthetic behavior estimation [5, 6, 7]. However, most of these works have focused on either a single locomotion task or jointly training multiple modes in the context of multi-locomotion scenarios. Further, prior works have framed learning from human demonstration as a standard supervised learning task, which assumes independence across time. This oversight can cascade predictive errors, deviating significantly from the original training distribution (refer also to prior work on imitation learning, such as [8, 9]). Moreover, the multifaceted dynamics of human gait, which can shift based on varying terrains, speeds, fatigue levels, or specific locomotion tasks, underscore the necessity for a model that fluidly adapts to these evolving conditions.
@@
To achieve this objective, we present a multitask, continually adaptive gait synergy approximation model tailored for various locomotion tasks, capable of adapting to evolving gait patterns and refining itself by integrating the impact of prediction errors. Central to our method is what we call the “multitask prospective rehearsal” (Figure 1) that prepares the model to anticipate and handle potential trajectories that may arise from its predictions, creating a seamless connection between continual adaptation and the integration of prediction errors for model refinement. Unlike conventional models that often rely on retrospective analysis of movement, our method is designed to prospectively imagine and synthesize potential future locomotion patterns. By synthesizing future states, the model is, in essence, creating its own new ’unseen’ data to practice on, which could lead to better generalization when encountering actual new data. The model refines its parameters incrementally as it updates itself with this enhanced data, negating the need for a simulator or new training data.
@@
We conduct a wide range of experiments in continual and joint learning settings with different model architectures and backbones on three real-world gait datasets, including our own patient (transtibial amputee) dataset, which we release for future use. We establish that our model outperforms many baseline models with multiple benchmarks.
Our study is the first to approach multi-gait adaptation in bionic prostheses as an error-aware multitask continual adaptation problem. Additionally, it is among the few that merge multitask adaptation with task-incremental learning in time series regression. Our methodology aims to produce set-point signals for bionic prostheses, achieving human-like movement fidelity and offering individuals with physical impairments the possibility of regaining near-normal functionality.
II Related Work
Learning-based Gait Behavior Models. In the field of predictive gait analytics, a variety of learning-based models have emerged, utilizing data from motion capture systems and wearable sensors to estimate a range of variables for diverse objectives. Ardestani et al. [10] modeled the medial knee contact force prediction using a feedforward artificial neural network (FFAN) based on motion capture data, offering a computational alternative to traditional techniques. Zhang et al. [11] harnessed wearable sensors combined with support vector regression to extract essential gait parameters efficiently, minimizing the need for extensive subject training and effectively addressing inter-subject variability. Concurrently, Rai et al. [12] employed recurrent neural networks (RNNs), integrating data from wearable sensors, to yield predictions of ankle angles from various ambulatory activities. Wang et al. [13] showcased the precision of a Gated Recurrent Unit (GRU) model for gait prediction from optical motion capture data. Livolsi et al. [14] introduced discrete wavelet transform methods for continuous gait phase extraction from hip encoder data. Kim et al. [15] reported the use of Long Short-Term Memory (LSTM) models to predict gait phases, analyzing short-term historical data on vertical loads and limb kinematics. The k-nearest neighbors (kNN) algorithm has been calibrated to work with pressure sensor inputs, enhancing the precision of walking phase prediction for prosthetic control systems [16]. To create more adaptive and responsive systems, Dey et al. [17] incorporated estimators that adjust to new training data within decision learner ensembles, catering to individual gait patterns. Waugh et al. [18] proposed a canonical dynamical system (CDS) model leveraging the Fourier series representation of gait. Yet, these solutions have generally been either constrained to single tasks, did not incorporate the effect of prediction errors, or have not tackled the problem of forgetting previously learned tasks when new ones are introduced.
@@ II Related Work
Multitask Adaptive Learning. A principal challenge in adaptive multitask learning is the phenomenon of catastrophic forgetting [19, 20], where newly acquired knowledge can cause the loss of previously learned information. To combat this, different continual learning strategies have been developed. Regularization-based methods like Elastic Weight Consolidation (EWC) [21] and Synaptic Intelligence (SI) [22] introduce regularization terms to guide the model towards a parameter space optimized for low error across both previous and new tasks. Evolving architectural methods such as Progressive Neural Networks (PNNs) [23] expand the model’s structure by introducing new, task-specific parameters to sustain performance across tasks. Rehearsal or replay methods, including Experience Replay (ER) and Gradient Episodic Memory (GEM) [24], maintain a sample of previous tasks’ data to be revisited during new task training, reinforcing the model’s exposure to the old data distribution. Rehearsal approaches have empirically proved to be the most effective ones among several approaches developed [25]. However, these techniques have a retrospective nature and are meant for treating data samples independently. In the context of bionic prosthetics, where predictive models interact with a dynamic environment, anticipating the impact of model predictions on subsequent data inputs is critical.
Model-based RL Model-based RL strategies generally focus on developing a representation of the environment by gathering data from interactions driven by a specific policy. This constructed model of the environment facilitates a level of planning, enabling the estimation of potential outcomes for future states [26, 27, 28, 29, 30]. Typically, these methods rely on a reward signal and online interactions with the environment. With the advent of offline reinforcement [31, 32] and model-based imitation learning [33, 34], access to the reward signals or online interactions can be bypassed [35]. As a result, it becomes feasible to simulate interactions and learn from these simulations, eliminating the necessity for real-time data acquisition. While model-based RL uses the world model for planning the agent’s actions based on its outcomes, we use it to generate rehearsal samples to fine-tune the prediction model to correct its own errors. Another major difference is that we use the world model in a multitask learning setting, whereas most of the model-based RL deals with single-task settings.
III Method
@@
Let $\mathbf{X}_{t}=\mathbf{x}_{t,1},\ldots,\mathbf{x}_{t,N_{t}}\in\mathbb{R}^{d}$ be a time series of sensor states (3D body and joint angles, angular velocities, and linear accelerations) during a locomotion task $t$ (such as level ground walking, ascending a ramp, or climbing stairs) and let $\mathbf{y}_{t}=y_{t,1},\ldots,y_{t,N_{t}}\in\mathbb{R}$ be the corresponding desired kinematic profiles (knee or ankle joint profiles) as demonstrated by a human. Our goal is a model $f$ which robustly predicts the correct kinematic profiles $y$ from input features $\mathbf{x}$, even under prediction errors in previous time steps, and adapts to new tasks $t$ while preserving performance on older tasks.
To realize such a model, we integrate principles from multitask adaptive learning with a prospective rehearsal approach inspired by model-based RL. Our model builds upon three core features, which we describe in turn: 1) multitask learning via a synthesis of shared learning mechanisms with task-specific adaptability, 2) continual adaptation with task rehearsal in the facet of an evolving architecture to avoid catastrophic forgetting, and 3) a prospective model that imagines and informs about potential deviations to provide robustness against prediction errors.
@@
III-A Multitask learning via generic and task-specific modules
@@
A core challenge in multitask learning is to exploit commonalities between the tasks while maintaining sufficient flexibility to account for the particularities of every single task. Our approach is to compose the gait model $f$ of two functions, namely a shared backbone $f^{s}$ across tasks, followed by task-specific layers $f_{t}$. The overall prediction for an input sample $\{\mathbf{x}_{t,i-T}...\mathbf{x}_{t,i}\}$ for task $t$ at time $i$ is $f(\mathbf{x}_{t,i-T}...\mathbf{x}_{t,i})=f_{t}[f^{s}(\mathbf{x}_{t,i-T}...\mathbf{x}_{t,i})]$. This duality enables the model to distill common features across all tasks and concurrently adapt to the unique aspects of each individual task. We model the shared backbone using temporal convolutions (TCN) [36, 7] due to its efficacy in handling temporal sequences. Task-specific layers are realized using a two-layer feed-forward network, ensuring a lightweight yet effective customization for each task.
III-B Continual adaptation via evolving architecture and rehearsal
@@
In (task) incremental adaptation, the model is trained incrementally with one task at a time. Thus, the model does not require data from all the tasks during training. When a new task $t$ is encountered, a new task-specific prediction layer $f_{t}$ is created and both $f_{t}$ and $f^{s}$ are trained using task-specific data. However, this method is prone to catastrophic forgetting [20, 19], that is, the model forgets the previously trained tasks when it learns a new task. To deal with the forgetting problem, we include samples from prior tasks in the training (i.e., a ’rehearsal’ [37]). However, beyond prior rehearsal approaches, we do not merely copy training data from prior tasks. Instead, we augment it, incorporating prediction errors from the current model $f$ using the following scheme.
III-C Prospective model
@@
A principal challenge in our scenario is that prediction errors at time step $k$ may negatively impact prediction performance at time step $k+1$, such that samples are not identically and independently distributed across time. In more detail, for the human demonstrations, we know that the sensor state $\mathbf{x}_{t,k}$ and desired kinematic profile $y_{t,k}$ is followed by the sensor state $\mathbf{x}_{t,k+1}$. However, if we predict a slightly different kinematic profile $\hat{y}_{t,k}=f(\mathbf{x}_{t,k-T}...\mathbf{x}_{t,k})\neq y_{t,k}$, we will also observe a different sensor state $\hat{\mathbf{x}}_{t,k+1}\neq\mathbf{x}_{t,k+1}$ in the next time step, pushing us away from the training data distribution. Such deviations can accumulate over time, which is a well-known phenomenon in imitation learning [8, 9]. In general, imitation learning is a helpful metaphor for our situation, as we also wish to mimic the demonstration of human experts in a setting where independence over time does not hold. However, past theory on imitation learning has usually assumed constant bounds on the error in every time step. Such constant bounds are not realistic in our scenario with a continuous space. Unfortunately, if we slightly relax the assumption of constant bounds to Lipschitz bounds, the deviation between desired time series and predicted time series can grow exponentially, as we show in the following, simple theorem.
Theorem III.1.
@@
Finally, assume that for all $k\in\{1,\ldots,N-1\}$, for all $x\in\mathcal{X}$, and for some $C\in\mathbb{R}$, the following Lipschitz condition holds: $d\Big{(}g\big{[}x,f(x)\big{]},x_{k+1}\Big{)}\leq C\cdot d(x,x_{k})$. Then, a) we obtain the exponential bound $d(x^{\prime}_{k},x_{k})\leq C^{k-1}\cdot d(x^{\prime}_{1},x_{1})$, and b) this bound is tight, i.e. there exist some time series $x_{1},\ldots,x_{N}$, some $f$ and $g$, as well as some $x^{\prime}_{1}$, such that the bound holds exactly.
@@
We obtain a) via induction from $d(x^{\prime}_{k+1},x_{k+1})=d\big{(}g[x^{\prime}_{k},f(x^{\prime}_{k})],x_{k+1}\big{)}\leq C\cdot d(x^{\prime}_{k},x_{k})$. Regarding b), consider the one-dimensional example $\mathcal{X}=\mathcal{Y}=\mathbb{R}$ with metric $d(x,x^{\prime})=|x-x^{\prime}|$, $x_{1}=\ldots=x_{N}=0$, $f(x)=\kappa\cdot x$ for some $\kappa>0$, and $g(x,y)=x+y$. Then, the Lipschitz condition is fulfilled because $d\big{(}g[x,f(x)],x_{k+1}\big{)}=|x+\kappa\cdot x-0|=|1+\kappa|\cdot|x-0|=(1+\kappa)\cdot d(x,x_{k})$, that is, $C=1+\kappa$. Further, for any $x^{\prime}_{1}\in\mathbb{R}$, we obtain $d(x^{\prime}_{k+1},x_{k+1})=|x^{\prime}_{k}+\kappa\cdot x^{\prime}_{k}-0|=(1+\kappa)\cdot|x^{\prime}_{k}-0|=C\cdot d(x^{\prime}_{k},x_{k})$, such that the tight bound follows via induction. ∎
@@
This observation motivates the key innovation of our work, namely a prospective rehearsal. More precisely, we train a prospective model $g$ to mimic the dynamics of the system with $g(\mathbf{x}_{k},{y}_{k})\approx\mathbf{x}_{k+1}$ and imagine the subsequent inputs resulting from the predictions of $f$. We then add data tuples $(\textbf{x}^{\prime}_{k+1},{y}_{k+1})$ with $\textbf{x}^{\prime}_{k+1}=g\big{(}\mathbf{x}_{k},f(\mathbf{x}_{k-T}...\mathbf{x}_{k})\big{)}$ to our training data. These additional training data tuples take the prediction errors of $f$ into account and thus help $f$ to counteract its own prediction errors. In our scenario, ${y}_{k+1}$ is the right target for the input $\textbf{x}^{\prime}_{k+1}$ because ${y}_{k+1}$ represents a target joint profile in a prosthetic limb. This target profile should remain the same in the same gait phase, even if $\textbf{x}^{\prime}_{k+1}$ slightly deviates from $\mathbf{x}_{k+1}$. In a general scenario, other targets might be required. Theoretically speaking, our aim is to achieve a constant bound $d\big{(}g[\mathbf{x}_{k},f(\mathbf{x}_{k-T},...,\mathbf{x}_{k})],\mathbf{x}_{k+1}\big{)}\leq\epsilon$ for small $\epsilon>0$, such that classical imitation learning theory applies [8, 9] and an exponential deviation over time is avoided. We do so by minimizing $d\big{(}g[\mathbf{x}_{k},{y}_{k}],\mathbf{x}_{k+1}\big{)}$ explicitly in Eq. (1).
III-D Summary of proposed approach
@@
$+\sum_{t^{\prime}=1}^{t-1}\sum_{k\in V_{t^{\prime}}}\left\lVert{y}_{t^{\prime},k}-f_{t^{\prime}}[f^{s}(\mathbf{x}_{t^{\prime},k},...,\mathbf{x}_{t^{\prime},k-T})]\right\rVert^{2}$
$+\left\lVert{y}_{t^{\prime},k}-f_{t^{\prime}}[f^{s}(\hat{\mathbf{x}}_{t^{\prime},k},...,\hat{\mathbf{x}}_{t^{\prime},k-T})]\right\rVert^{2}$

### v3

# Continual Learning from Simulated Interactions via Multitask Prospective Rehearsal for Bionic Limb Behavior Modeling

@@ Abstract
Lower limb amputations and neuromuscular impairments severely restrict mobility, necessitating advancements beyond conventional prosthetics. While motorized bionic limbs show promise, their effectiveness depends on replicating the dynamic coordination of human movement across diverse environments. In this paper, we introduce a model for human behavior in the context of bionic prosthesis control. Our approach leverages human locomotion demonstrations to learn the synergistic coupling of the lower limbs, enabling the prediction of the kinematic behavior of a missing limb during tasks such as walking, climbing inclines, and stairs. We propose a multitasking, continually adaptive model that anticipates and refines movements over time. At the core of our method is a technique called multitask prospective rehearsal, that anticipates and synthesizes future movements based on the previous prediction and employs a corrective mechanism for subsequent predictions. Our evolving architecture merges lightweight, task-specific modules on a shared backbone, ensuring both specificity and scalability. We validate our model through experiments on real-world human gait datasets, including transtibial amputees, across a wide range of locomotion tasks. Results demonstrate that our approach consistently outperforms baseline models, particularly in scenarios with distributional shifts, adversarial perturbations, and noise.
Keywords: Human Behavior Modeling, Bionics, Multitask Rehearsal, World Model, Human-machine Interaction
1 Introduction
@@
Lower limb amputation and neuromuscular disorders detrimentally affect natural locomotion, compromising individuals’ quality of life (Windrich et al., 2016). Those afflicted often depend on assistive technologies like prosthetics or orthoses to regain daily mobility (Windrich et al., 2016). However, conventional passive prosthetics often fall short in replicating natural gait during diverse activities, from basic transitions like standing from a seated position to more dynamic actions like running or navigating slopes (Windrich et al., 2016). The advent of powered prosthetics, equipped with integrated motors, promises a more naturalistic gait. Nonetheless, this advancement requires a model capable of effectively approximating the complexities of human gait synergy to accurately estimate the necessary motor commands. While finite-state machines are adequate for rudimentary scenarios (Lawson, 2014; Chen et al., 2015), their construction becomes infeasible for finer gait phase resolution and multiple locomotion tasks (Lawson et al., 2013). An emerging solution lies in training models on human demonstrations to intuitively infer target limb motion (Dey et al., 2019; 2020a; 2021b). This methodology not only replicates gait’s inherent fluidity but also facilitates seamless transitions across varying gaits, eliminating the need for rigid rules or heuristics. Furthermore, such learning-based gait models can be robustified by training on a large and diverse set of human demonstrations Dey et al. (2020b). To achieve personalization, these models could incorporate references that exhibit anthropomorphic features similar to those of the physically impaired subject Dey & Nair (2024).
@@
Recent studies attest to the efficacy of learning-based models in gait and prosthetic behavior estimation (Dhir et al., 2018; Zabre-Gonzalez et al., 2021; Fang et al., 2020; Dey & Schilling, 2022a). However, most of these works have focused on either a single locomotion task or jointly training multiple modes in the context of multi-locomotion scenarios. Furthermore, prior research has approached multitask learning from human demonstration as a standard supervised learning task, which assumes independence across time. This oversight can cascade predictive errors, deviating significantly from the original training distribution (Kumar et al., 2022; Ross & Bagnell, 2010). Some approaches (Ross et al., 2011; Judah et al., 2014; Kelly et al., 2019) have addressed the issue of state distribution shift by using a trained model to collect additional data and label them as the model encounters new states, refining the model iteratively. However, these studies require having access to a simulator, environment, or deployment of models in the real world for data collection to estimate the expert action for the induced state. This may not be feasible or highly costly for multiple safety-critical scenarios, e.g., human-centered robotics that might entail risks of imbalance and fall. Moreover, these approaches have largely focused on single-task settings rather than continual multitask adaptation. The multifaceted dynamics of human gait, which can shift based on varying terrains, speeds, fatigue levels, or specific locomotion tasks, underscore the necessity for a model that fluidly adapts to these evolving conditions.
@@
To achieve this objective, we present a multitask, continually adaptive gait synergy approximation model tailored for various locomotion tasks, capable of adapting to evolving gait patterns and refining itself by integrating the impact of prediction errors. Central to our method is what we call the multitask prospective rehearsal (Figure 1) that prepares the model to anticipate and handle potential trajectories that may arise from its predictions, creating a seamless connection between continual adaptation and the integration of prediction errors for model refinement. Unlike conventional models that often rely on retrospective analysis of movement, our method is designed to prospectively imagine and synthesize potential future locomotion patterns. By synthesizing future states, the model is, in essence, creating its own new ’unseen’ data to practice on, which could lead to better generalization when encountering actual new data. The model refines its parameters incrementally as it updates itself with this enhanced data, negating the need for a simulator or new training data.
@@
We conduct a wide range of experiments in continual and joint learning settings with different model architectures and backbones on three real-world gait datasets, including our own patient (transtibial amputee) dataset. We establish that our model outperforms many baseline models with multiple benchmarks. Our study is the first to approach multi-gait adaptation in bionic prostheses as an error-aware multitask continual adaptation problem.
2 Related Work
Prosthetic behavior models. Conventional prosthetic behavior models primarily rely on state machines that deterministically generate control commands based on finite states within a gait cycle. These models segment the continuous gait cycle into discrete states (Lawson, 2014; Chen et al., 2015; Culver et al., 2023), which fails to accurately replicate the smooth and continuous movement patterns typical of human walking. To better model the fluidity of the gait cycle, learning-based regression methods (Dey et al., 2020b; 2021a; Dhir et al., 2018; Fang et al., 2020; Zabre-Gonzalez et al., 2021; Dey & Schilling, 2022b) have been utilized that directly predict the joint motion commands from the sensor states of residual body movement. However, these models generally adopt a straightforward reactive strategy, learning the correlation between the residual states of the body and the limb’s movements. Hence, these models often oversimplify the complex nature of human motor abilities, which are inherently adaptable to various tasks, diverse environments, and evolving gait patterns.
@@ 2 Related Work
Multitask adaptive learning. A principal challenge in multitask adaptive learning is the phenomenon of catastrophic forgetting (McCloskey & Cohen, 1989; Robins, 1995; French, 1999), where newly acquired knowledge can cause the loss of previously learned information. To combat this, continual learning strategies, such as regularization-based methods, architectural strategies, and experience replay (ER)(Kirkpatrick et al., 2017; Zenke et al., 2017; Smith et al., 2023) have been developed. Regularization-based methods like Elastic Weight Consolidation (EWC) (Kirkpatrick et al., 2017) and Synaptic Intelligence (SI) (Zenke et al., 2017) introduce regularization terms to guide the model towards a parameter space optimized for low error across both previous and new tasks. Architectural methods such as Progressive Neural Networks (PNNs) (Rusu et al., 2016) expand the model’s structure by introducing task-specific parameters to sustain performance across tasks. Rehearsal or replay methods, including Experience Replay (ER) (Robins, 1995; Luo & Li, 2020; Li et al., 2023) and Gradient Episodic Memory (GEM) (Lopez-Paz & Ranzato, 2017), maintain a sample of previous tasks to be revisited during new task training, reinforcing the model’s exposure to the old data distribution. Rehearsal approaches have empirically proved to be the most effective among several approaches developed (Merlin et al., 2022). However, these techniques have a retrospective nature and are meant for treating data samples independently. In the context of bionic prosthetics, where predictive models interact with a dynamic environment, anticipating the impact of predictions on subsequent data inputs is critical.
Model-based reinforcement learning (RL). Model-based RL focuses on developing a representation of the environment by gathering data from interactions driven by a specific policy. This constructed model of the environment facilitates a level of planning, enabling the estimation of potential outcomes for future states (Mnih et al., 2013; Silver et al., 2016; Racanière et al., 2017; Nagabandi et al., 2018; Chua et al., 2018). Typically, these methods rely on a reward signal and online interactions with the environment. With the advent of offline reinforcement (Zhou et al., 2021; Yu et al., 2020) and model-based imitation learning (IL) (Englert et al., 2013; Kidambi et al., 2021), access to the reward signals or online interactions can be bypassed (Hu et al., 2022). Both model-based RL and IL, by internalizing environmental dynamics, enable an agent to forecast outcomes and make informed decisions. Our approach is inspired by model-based RL; however, we use it as a rehearsal strategy within a multitask continual learning setting to mitigate forgetting, adapt to evolving tasks, and limit compounding errors, all without needing access to a reward signal or online interactions.
3 Method
@@
Let $\mathbf{X}_{t}=\mathbf{x}_{t,1},\ldots,\mathbf{x}_{t,N_{t}}\in\mathbb{R}^{d}$ be a time series of sensor states (3D body and joint angles, angular velocities, and linear accelerations) during a locomotion task $t$ (such as level ground walking, ascending a ramp, or climbing stairs) and let $\mathbf{y}_{t}=y_{t,1},\ldots,y_{t,N_{t}}\in\mathbb{R}$ be the corresponding desired kinematic profiles (knee or ankle joint profiles) as demonstrated by a human. Our goal is a model $f$ which robustly predicts the correct kinematic profiles $y$ from input features $\mathbf{x}$, even under prediction errors in previous time steps, and adapts to new tasks $t$ while preserving performance on older tasks. To realize such a model, we integrate principles from multitask adaptive learning with a prospective rehearsal approach inspired by model-based RL. Our model builds upon three core features, which we describe in turn: 1) multitask learning via a synthesis of shared learning mechanisms with task-specific adaptability, 2) continual adaptation with task rehearsal in the facet of an evolving architecture to avoid catastrophic forgetting, and 3) a prospective model, that imagines and informs about potential deviations to provide robustness against prediction errors.
@@
Multitask learning via generic and task-specific modules. A core challenge in multitask learning is to exploit commonalities between the tasks while maintaining flexibility to account for the particularities of every single task. Our approach is to compose the gait model $f$ of two functions: a shared backbone $f^{s}$ across tasks, followed by task-specific layers $f_{t}$. The overall prediction for an input sample $\{\mathbf{x}_{t,i-T}...\mathbf{x}_{t,i}\}$ for task $t$ at time $i$ is $f(\mathbf{x}_{t,i-T}...\mathbf{x}_{t,i})=f_{t}[f^{s}(\mathbf{x}_{t,i-T}...\mathbf{x}_{t,i})]$. This duality enables the model to distill common features across all tasks and concurrently adapt to the unique aspects of individual tasks. We employ temporal convolutions (Oord et al., 2016; Fang et al., 2020) to model the shared backbone due to their effectiveness in handling temporal sequences. To tailor the model to specific tasks, we implement task-specific layers using a two-layer feed-forward network, providing lightweight and efficient customization for each task.
@@
Continual learning via evolving architecture and rehearsal. In (task) incremental adaptation, the model is trained incrementally with one task at a time. Thus, the model does not require data from all the tasks during training. When a new task $t$ is encountered, a new task-specific prediction layer $f_{t}$ is created and both $f_{t}$ and $f^{s}$ are trained using task-specific data. However, this method is prone to catastrophic forgetting (French, 1999; McCloskey & Cohen, 1989), that is, the model forgets the previously trained tasks when it learns a new task. To deal with the forgetting problem, we include samples from prior tasks in the training (i.e., a ’rehearsal’ (van de Ven et al., 2020)). However, beyond prior rehearsal approaches, we do not merely copy training data from prior tasks. Instead, we augment it, incorporating prediction errors from the current model $f$ using the following scheme.
@@
Prospective model. A principal challenge in our scenario is that prediction errors at time step $k$ may negatively impact prediction performance at time step $k+1$, such that samples are not identically and independently distributed across time. In more detail, for the human demonstrations, we know that the sensor state $\mathbf{x}_{t,k}$ and desired kinematic profile $y_{t,k}$ is followed by the sensor state $\mathbf{x}_{t,k+1}$. However, if we predict a slightly different kinematic profile $\hat{y}_{t,k}=f(\mathbf{x}_{t,k-T}...\mathbf{x}_{t,k})\neq y_{t,k}$, we will also observe a different sensor state $\hat{\mathbf{x}}_{t,k+1}\neq\mathbf{x}_{t,k+1}$ in the next time step, pushing us away from the training data distribution. Such deviations can accumulate over time, which is a well-known phenomenon in imitation learning (Kumar et al., 2022; Ross & Bagnell, 2010). In general, imitation learning is a helpful metaphor for our situation, as we also wish to mimic the demonstration of human experts in a setting where independence over time does not hold. However, past theory on imitation learning has usually assumed constant bounds on the error in every time step. Such constant bounds are not realistic in our scenario with a continuous space. Unfortunately, if we slightly relax the assumption of constant bounds to Lipschitz bounds, the deviation between desired time series and predicted time series can grow exponentially, as we show in the following, simple theorem.
Theorem 1.
@@
Finally, assume that for all $k\in\{1,\ldots,N-1\}$, for all $x\in\mathcal{X}$, and for some $C\in\mathbb{R}$, the following Lipschitz condition holds: $d\Big{(}g\big{[}x,f(x)\big{]},x_{k+1}\Big{)}\leq C\cdot d(x,x_{k})$. Then, a) we obtain the exponential bound $d(x^{\prime}_{k},x_{k})\leq C^{k-1}\cdot d(x^{\prime}_{1},x_{1})$, and b) this bound is tight, i.e. there exist at least one combination of time series $x_{1},\ldots,x_{N}$, some $f$ and $g$, as well as some $x^{\prime}_{1}$, such that the bound holds exactly.
@@
We obtain a) via induction: The base case is $d(x^{\prime}_{1},x_{1})\leq C^{0}\cdot d(x^{\prime}_{1},x_{1})$, which is trivially fulfilled. Now, assume that the claim holds for $k$. Then, for $k+1$, we obtain $d(x^{\prime}_{k+1},x_{k+1})=d\big{(}g[x^{\prime}_{k},f(x^{\prime}_{k})],x_{k+1}\big{)}\leq C\cdot d(x^{\prime}_{k},x_{k})$ due to the definition of $x^{\prime}_{k+1}$ and the Lipschitz assumption made in the theorem. According to the induction hypothesis, the right-hand-side is bounded by $C\cdot C^{k-1}\cdot d(x^{\prime}_{1},x_{1})=C^{k}\cdot d(x^{\prime}_{1},x_{1})$, which concludes the induction.
b) First, note that it is sufficient to show one specific example to prove the tightness of the bound as we only claim that there exists at least one example for which the bound holds exactly. Then, assuming the setup of the example as given in the proof, we obtain the following induction. The base case is $d(x^{\prime}_{1},x_{1})\leq C^{0}\cdot d(x^{\prime}_{1},x_{1})$, which is trivially fulfilled. Now, assume the claim holds for $k$. Then, for $k+1$, we obtain $d(x^{\prime}_{k+1},x_{k+1})=|x^{\prime}_{k}+\kappa\cdot x^{\prime}_{k}-0|=(1+\kappa)\cdot|x^{\prime}_{k}|=C\cdot|x^{\prime}_{k}-0|=C\cdot|x^{\prime}_{k}-x_{k}|=C\cdot d(x^{\prime}_{k},x_{k})$, utilizing the definitions of $x_{1}=\ldots=x_{N}=0,d(x,y)=|x-y|$, and $C=1+\kappa$. According to the induction hypothesis, the right-hand-side is equal to $C\cdot C^{k-1}\cdot d(x^{\prime}_{1},x_{1})=C^{k}\cdot d(x^{\prime}_{1},x_{1})$, which concludes the induction. Accordingly, for this example, the equality is always fulfilled exactly, implying that the bound is tight. ∎
@@
This observation motivates the key innovation of our work, namely a prospective rehearsal. More precisely, we train a prospective model $g$ to mimic the dynamics of the system with $g(\mathbf{x}_{k},{y}_{k})\approx\mathbf{x}_{k+1}$ and imagine the subsequent inputs resulting from the predictions of $f$. We then add data tuples $(\textbf{x}^{\prime}_{k+1},{y}_{k+1})$ with $\textbf{x}^{\prime}_{k+1}=g\big{(}\mathbf{x}_{k},f(\mathbf{x}_{k-T}...\mathbf{x}_{k})\big{)}$ to our training data. These additional training data tuples take the prediction errors of $f$ into account and thus help $f$ to counteract its own prediction errors. In our scenario, ${y}_{k+1}$ is the right target for the input $\textbf{x}^{\prime}_{k+1}$ because ${y}_{k+1}$ represents a target joint profile in a prosthetic limb. This target profile should remain the same in the same gait phase, even if $\textbf{x}^{\prime}_{k+1}$ slightly deviates from $\mathbf{x}_{k+1}$. In a general scenario, other targets might be required. Theoretically speaking, our aim is to achieve a constant bound $d\big{(}g[\mathbf{x}_{k},f(\mathbf{x}_{k-T},...,\mathbf{x}_{k})],\mathbf{x}_{k+1}\big{)}\leq\epsilon$ for small $\epsilon>0$, such that classical imitation learning theory applies (Kumar et al., 2022; Ross & Bagnell, 2010) and an exponential deviation over time is avoided. We do so by minimizing $d\big{(}g[\mathbf{x}_{k},{y}_{k}],\mathbf{x}_{k+1}\big{)}$ explicitly in Eq. equation 1.
Summary of proposed approach. Our overall approach is shown in Algorithm 1 and Figures 1 and 2. We add new tasks $t$ one by one and split the training data for task $t$ into a training and a validation set. We train a newly evolved task-specific layer $f_{t}$ and the shared backbone $f^{s}$ to minimize the distance between $f_{t}[f^{s}(\mathbf{x}_{t,k-T}...\mathbf{x}_{t,k})]$ and ${y}_{t,k}$, and, at the same time, we train the shared backbone $f^{s}$ and the old task layers $f_{t^{\prime}}$ to minimize the distance between $f_{t^{\prime}}[f^{s}(\mathbf{x}_{t^{\prime},k-T}...\mathbf{x}_{t^{\prime},k-1},\hat{\mathbf{x}}_{t^{\prime},k})]$ and ${y}_{t^{\prime},k}$ for the data in the rehearsal buffer. The rehearsal buffer contains both the original validation samples $\mathbf{x}_{t^{\prime},k}$, and the prospective samples $\hat{\mathbf{x}}_{t^{\prime},k}$ for a subsample $V_{t^{\prime}}$ of the time steps $k$ in the validation set. The prospective samples are generated via a task-specific, prospective model $g_{t^{\prime}}$ as $\hat{\mathbf{x}}_{t^{\prime},k}=g_{t^{\prime}}\big{(}\mathbf{x}_{t^{\prime},k-1},f_{t^{\prime}}[f^{s}(\mathbf{x}_{t^{\prime},k-T-1}...\mathbf{x}_{t^{\prime},k-1})]\big{)}$.
@@
$+\sum_{t^{\prime}=1}^{t-1}\sum_{k\in V_{t^{\prime}}}\left\lVert{y}_{t^{\prime},k}-f_{t^{\prime}}[f^{s}(\mathbf{x}_{t^{\prime},k},...,\mathbf{x}_{t^{\prime},k-T})]\right\rVert^{2}+\left\lVert{y}_{t^{\prime},k}-f_{t^{\prime}}[f^{s}(\hat{\mathbf{x}}_{t^{\prime},k},...,\hat{\mathbf{x}}_{t^{\prime},k-T})]\right\rVert^{2}$